# A Comparison-Relationship-Surrogate Evolutionary Algorithm for Multi-Objective Optimization

### v1

@@ Abstract
Evolutionary algorithms often struggle to find high-quality solutions to multi-objective optimization problems on a limited budget of function evaluations (here, a few hundred). A promising direction to improve the efficiency of these methods is to augment the objective functions with a data-driven surrogate model. These “surrogate-assisted” optimization algorithms can achieve better solutions than conventional algorithms for the same number of function evaluations on a wide variety of test problems. In this work, we continue to explore the area of surrogate-assisted multi-objective optimization by introducing and testing an algorithm driven by a new type of surrogate model: a comparison-relationship-surrogate model. This model predicts the truth values of the comparison operator evaluated on the objective functions for two candidate solutions. These predictions can be used to infer the domination relationships that power the non-dominated sorting mechanism used by many multi-objective genetic algorithms to select fit individuals. Several numerical experiments are performed on this algorithm using well-known test suites plus a real-world problem from the field of accelerator physics. Statistical analysis of the results demonstrates that the new algorithm can, on average, achieve better-converged solutions to many medium-scale, biobjective problems than existing state-of-the-art methods for a limited budget of function evaluations.
@@
Surrogate-Assisted Evolutionary Algorithm (SAEA), Expensive Optimization, Medium-Scale Multi-objective Optimization, Comparison Relationship Surrogate, Pairwise Comparison.
journal: Swarm and Evolutionary Computation
@@ 1 Introduction
A popular approach to increasing the sample efficiency of multi-objective genetic algorithms is to introduce a surrogate model: a fast-to-evaluate model of the objectives based on already evaluated data. These surrogate models (also called meta-models) can act as a stand-in for the expensive objective function evaluations in existing optimization algorithms. For example, in K-RVEA [11], a model of the objective functions is generated from an initially evaluated set of individuals and the optimization is performed against the model using a reference-vector-based evolutionary algorithm. The model is then periodically updated with real calls to the objective functions for candidate solutions. For a review of surrogate-assisted evolutionary algorithms (SAEAs), readers are pointed to [9, 12, 13].
@@
    A surrogate model of the comparison relationships between the objective functions of two individuals is proposed for solving expensive multi-objective optimization problems. This model is implemented using a neural network with an architecture designed to reflect the symmetry properties of the comparison operator.
@@
    A method of reconstructing the ranking of individuals from their approximate comparisons (which may not satisfy the transitive property) is explored.
@@
The rest of this article is dedicated to exploring the use of a comparison-relationship-surrogate model to improve the sample efficiency of multi-objective genetic algorithms. In Sec. 2, we review the existing literature surrounding surrogate-assisted evolutionary algorithms (SAEAs) which are powered with classification-based models and the state of comparison relationship models for optimization. We describe the new SAEA that uses a model of the comparison relationships in Sec. 3. The construction of the model using a neural network is also described and architectural / training details related to the symmetries of the comparison operator are discussed. In Sec. 4, we report on numerical experiments that compare the new algorithm to a diverse set of existing SAEAs. These experiments are performed on several well-known suites of test problems. Results from the algorithm on an example real-world optimization problem are also given in Sec. 5. We conclude the paper in Sec. 6 and give suggestions for future areas of study related to the comparison-relationship-surrogate optimization algorithm.
@@ 2.1 Classification-Based SAEAs
Surrogate assisted evolutionary algorithms driven by classification models have shown some success in improving the efficiency with which expensive optimization problems can be solved on a budget. In early work, the algorithm Pareto-SVM [14] was introduced which uses a support vector machine to classify regions of decision space into those which map to points in objective space that are part of the dominated set, the current Pareto front, or unexplored regions. A variant of the algorithm [15] followed shortly after and introduced a new classification model that accepts pairs of individuals and approximates their domination relationship directly. The themes of classifying either regions of decision space or the relationships between pairs of individuals can be seen throughout later work and we use them to help organize the remainder of this section.
@@
We begin by taking a look at some of the algorithms that use classification models that accept a single point in decision space as their input. In CPS-MOEA [16], the individuals in the training set are labeled into equally sized positive and negative classes using a combination of nondomination rank and crowding distance. A k-nearest neighbor classifier is then trained on the data and used to prefilter candidate solutions before evaluation. The algorithm MOEA/D-SVM [17] explored using a scalarization function on the objectives to split the population into positive and negative classes. In one proposed algorithm [18], a “one-class classification model” avoided the problem of splitting the data by only considering positive training examples. CSA-MOEA [19] explored using a decision tree to sort regions of decision space into those which were and were not promising to find candidates in. New infill strategies were explored for choosing which candidates to evaluate. Multiple surrogate models were applied to solve “local” problems in a decomposition-based optimization algorithm to mitigate overfitting in MCEA/D [20]. The algorithm CR-SAEA [21] augments the multiple local classification models approach with a global regression model to help guide its decomposition-based genetic algorithm.
@@
Modeling the relationships between individuals (eg domination) is also an important theme in multi-objective optimization algorithms aided by classification models. The concept of modeling the Pareto-dominance relationship was explored further in some work [22] which compared ten off-the-shelf classification models for use in this task. The algorithm CSEA [23] was introduced which uses a surrogate model to estimate domination between candidate solutions and selected already evaluated reference solutions. The algorithm also uses a measure of uncertainty in the predictions to choose which candidates to evaluate. Two surrogate models: one of the Pareto-domination relationships and a second of the $\theta$-domination relationships between candidates are combined in the algorithm $\theta$-DEA-DP [24]. In REMO [25], the relationships between candidate solutions are modeled with a neural network classifier trained on pairs of individuals which are grouped into equal sized sets based on domination. A related algorithm, named CREMO [26], also explored relationship learning using a convolutional neural network on the decision variables. In FAEA-SDR [27], the strengthened dominance relationship between individuals is modeled instead. Pairwise comparisons of the scalarized objective functions are learned and used in the optimization algorithm PC-SAEA [28]. The concept of domination modeling has also been applied to neural architecture search with promising results [29].
@@
Some work has been put into modeling the comparison relationships between the objective functions directly, but this has remained restricted to the case of single-objective optimization or constraint satisfaction problems. One such study [30] showed an algorithm that ranked individuals using a support vector machine could outperform evolutionary algorithms aided with regression models on some benchmark problems. Different classification models were compared for ranking candidates in single-objective optimization problems in [31]. A wide variety of pairwise classification models and regression models were compared against each other on benchmark problems for single-objective optimization in [32] with some evidence for the success of the classification approach. A specially designed symmetric neural network architecture was created for the task of modeling the comparison operator between multiple objectives in [33] and applied to solve constraint satisfaction problems in integrated circuit design. While these results demonstrate the potential of a comparison-based model, they did not address the unique challenges of integrating one with a multi-objective optimization algorithm.
@@
2 $P_{t}\leftarrow$ initial set of $N_{i}$ individuals, Latin hypercube
3 $P\leftarrow$ $N$ individuals selected from $P_{t}$ as in NSGA-II
@@ 11             $P_{m}\leftarrow P_{m}\cup c_{m}$
12             $P_{m}\leftarrow$ select $N$ individuals in $P_{m}$ using model
@@ 14             $w\leftarrow w+1$
      /* Evaluate individuals */
18       $c\leftarrow$ evaluate $\mu$ random individuals from $P_{m}$
@@ 21       $\theta\leftarrow$ train model on $P_{t}$
      /* Select fit individuals */
@@ 23       $P\leftarrow P\cup c$
24       $P\leftarrow$ select $N$ individuals from $P$ as in NSGA-II
@@
Figure 1: A diagram of the how the CRSEA algorithm (Alg. 1) works. An initial population is generated and the model trained. Then, the model is used in the place of the objective functions in an evolutionary algorithm for some number of generations ($w_{\text{max}}$). Several ($\mu$) randomly chosen individuals from the resulting candidate solutions are evaluated and used to retrain the model. The process is repeated until the budget of function evaluations ($\text{FE}_{max}$) is exhausted.
@@ 2.2 Motivation
In addition to the invariance properties of the model, a classification problem is always “simpler” than the equivalent regression problem in the sense that it is solved in the process of solving the regression problem, but does not solve the regression problem itself. That is, a model of the objective function values also gives the classifications whether they predict domination, comparisons, or something else. However, knowledge of what class an individual lies in is not enough information to determine the values of the objective functions. It may be that this “simpler” problem is easier to solve.
The task of learning the comparison relationships between the objectives also forms a naturally balanced classification problem. In a model of the domination relationships, for example, as the algorithm drifts closer to the Pareto front most new examples are dominated by the existing population. The comparisons between objective functions, on the other hand, can always be manipulated to have an equal number of positive and negative examples by swapping the order of the two individuals being compared.
With these arguments and the previously discussed success of comparison relationship models in solving single-objective optimization problems, we are motivated to use a similar model to drive a genetic algorithm in solving the multi-objective case.
@@ 3 A Comparison-Relationship-Surrogate multi-objective Genetic Algorithm
With an accurate enough model of the comparison relationships (Eq. 1), the expense of evaluating the objective functions in an evolutionary algorithm could be replaced completely with a cheaper call to the model. With this intuition, we propose the algorithm described in Alg. 1. The model is first initialized using training points that are drawn from the decision space with Latin hypercube sampling [37] and evaluated using the real objective functions. Candidate solutions are found from the model using an evolutionary algorithm. To combat the model’s loss of accuracy when it is evaluated far from the points it was trained on, the model is periodically updated using a small number of candidate solutions for which the real objective function values are computed. This process (solving the optimization problem and updating the model) is repeated until the budget of function evaluations is exhausted. A diagram of the algorithm is shown in Fig. 1.
3.1 Calculating an Approximate Pareto Front From the Model
@@
With this restriction in mind, we propose finding optimal solutions from the model using NSGA-II [38]. Briefly summarizing the algorithm, in each generation, a set of children is produced with mutation, crossover, and selection operators (in this paper, polynomial mutation, simulated binary crossover, and tournament selection [39]). After evaluating the objective functions, the population is sorted by increasing rank of non-domination. Within those ranks, the “crowding distance” metric is used to break ties in order to promote diversity. Individuals are selected from the newly sorted list with the smallest rank, largest crowding distance individuals being the most fit. We suggest running the algorithm using only information from the surrogate model.
Although NSGA-II is well suited for use with the comparison-relationship model, there are still a few challenges that come with it. Each is described in detail in the following subsections and we propose ways of mitigating them.
@@ 3.1.1 Violation of the Transitive Property
Now that we can use the model to compute an approximate version of the domination relationship, we run into a different type of problem using it with non-dominated sorting algorithms. The issue is that the model is only an approximation of the comparison operator and it isn’t guaranteed to respect the transitive property,
$(a<b)\wedge(b<c)\implies a<c,$ (2)
where $\wedge$ is the logical AND operation. Besides causing issues with accuracy, depending on how badly the model violates this condition, the fast non-dominated sorting algorithm described in [38] can produce nonsensical results.
We propose “cleaning” the predictions of the comparison model so that they respect the transitive property. For all individuals $x_{i}$ in $\mathbb{P}$, the population being sorted, we can calculate the noisy predictions from the comparison model. The predictions take the form of a probability that the comparison is true, $\mathcal{P}(f(x_{i})<f(x_{j}))$, where $f(x)$ is the objective function (a single objective for simplicity). To clean the predictions, we calculate the value,
@@
from Ref. [40]. The diagonal of the matrix in the sum is defined as $\mathcal{P}(f(x_{i})<f(x_{i}))=0$. The quantity $\tau_{i}$ can be interpreted as the probability of the objective function of individual $i$ being smaller than that of an individual chosen at random from $\mathbb{P}$. The comparison predictions defined by $f(x_{i})<f(x_{j}):=\tau_{i}>\tau_{j}$ now satisfy the transitive property. This cleaning procedure is repeated for all objective functions and the domination relationships can be calculated from there, allowing the fast non-dominated sorting algorithm to be run using only the comparison model.

### v2

@@ Abstract
Evolutionary algorithms often struggle to find well converged (e.g small inverted generational distance on test problems) solutions to multi-objective optimization problems on a limited budget of function evaluations (here, a few hundred). The family of surrogate-assisted evolutionary algorithms (SAEAs) offers a potential solution to this shortcoming through the use of data driven models which augment evaluations of the objective functions. A surrogate model which has shown promise in single-objective optimization is to predict the “comparison relationship” between pairs of solutions (i.e. who’s objective function is smaller). In this paper, we investigate the performance of this model on multi-objective optimization problems. First, we propose a new algorithm “CRSEA” which uses the comparison-relationship model. Numerical experiments are then performed with the DTLZ and WFG test suites plus a real-world problem from the field of accelerator physics. We find that CRSEA finds better converged solutions than the tested SAEAs on many of the medium-scale, biobjective problems chosen from the WFG suite suggesting the “comparison-relationship surrogate” as a promising tool for improving the efficiency of multi-objective optimization algorithms.
@@
Surrogate-Assisted Evolutionary Algorithm (SAEA), Expensive Optimization, Medium-Scale Multi-objective Optimization, Comparison-Relationship Surrogate, Pairwise Comparison.
@@ 1 Introduction
A popular approach to increasing the sample efficiency of multi-objective genetic algorithms is to introduce a surrogate model: a fast-to-evaluate model of the objectives based on already evaluated data. These surrogate models (also called meta-models) can act as a stand-in for the expensive objective function evaluations in existing optimization algorithms. For example, in K-RVEA [11], a model of the objective functions is generated from an initially evaluated set of solutions and the optimization is performed against the model using a reference-vector-based evolutionary algorithm. The model is then periodically updated with real calls to the objective functions for candidate solutions. For a review of surrogate-assisted evolutionary algorithms (SAEAs), readers are pointed to [9, 12, 13].
@@
    A surrogate model of the comparison relationships between the objective functions of two solutions is proposed for solving expensive multi-objective optimization problems. This model is implemented using a neural network with an architecture designed to reflect the symmetry properties of the comparison operator.
@@
    A method of reconstructing the ranking of solutions from their approximate comparisons (which may not satisfy the transitive property) is explored.
@@
The rest of this article is dedicated to exploring the use of a comparison-relationship-surrogate model to improve the sample efficiency of multi-objective genetic algorithms. In Sec. 2, we review the existing literature surrounding surrogate-assisted evolutionary algorithms (SAEAs) which are powered with classification-based models and the state of comparison-relationship models for optimization. We describe the new SAEA that uses a model of the comparison relationships in Sec. 3. The construction of the model using a neural network is also described and architectural / training details related to the symmetries of the comparison operator are discussed. In Sec. 4, we report on numerical experiments that compare the new algorithm to a diverse set of existing SAEAs. These experiments are performed on several well-known suites of test problems. Results from the algorithm on an example real-world optimization problem are also given in Sec. 5. We conclude the paper in Sec. 6 and give suggestions for future areas of study related to the comparison-relationship-surrogate optimization algorithm.
@@ 2 Related Work
2.1 Regression-Based SAEAs
One approach to building a data-driven model of the already evaluated solutions is to treat the situation as a regression problem. We can attempt to predict the values of the objective functions or metrics related to them using the decision variables.
The algorithm ParEGO [14], for example, uses a Gaussian process model of a scalarization function of the objectives. New points are chosen to maximize the expected improvement of the solution in a manner similar to the kriging method [15]. SMS-EGO extended the idea to maximize improvement in the hypervolume (S-metric) instead [16]. In K-RVEA, the regression model is used directly in place of the expensive objective function evaluations in a reference-vector guided evolutionary algorithm [11]. Multi-objective optimization problems are first decomposed into single objective sub-problems in the algorithm MOEA/D-EGO [17]. A Gaussian process is used to model each of the scalar sub-problems which are then solved simultaneously to suggest new candidate solutions to evaluate.
@@ 2.2 Classification-Based SAEAs
Surrogate assisted evolutionary algorithms driven by classification models have shown some success in improving the efficiency with which expensive optimization problems can be solved on a budget. In early work, the algorithm Pareto-SVM [18] was introduced which uses a support vector machine to classify regions of decision space into those which map to points in objective space that are part of the dominated set, the current Pareto front, or unexplored regions. A variant of the algorithm [19] followed shortly after and introduced a new classification model that accepts pairs of solutions and approximates their domination relationship directly. The themes of classifying either regions of decision space or the relationships between pairs of solutions can be seen throughout later work and we use them to help organize the remainder of this section.
2.2.1 Decision Space Classification
@@
We begin by taking a look at some of the algorithms that use classification models that accept a single point in decision space as their input. In CPS-MOEA [20], the solutions in the training set are labeled into equally sized positive and negative classes using a combination of non-domination rank and crowding distance. A k-nearest neighbor classifier is then trained on the data and used to pre-filter candidate solutions before evaluation. The algorithm MOEA/D-SVM [21] explored using a scalarization function on the objectives to split the population into positive and negative classes. In one proposed algorithm [22], a “one-class classification model” avoided the problem of splitting the data by only considering positive training examples. CSA-MOEA [23] explored using a decision tree to sort regions of decision space into those which were and were not promising to find candidates in. New infill strategies were explored for choosing which candidates to evaluate. Multiple surrogate models were applied to solve “local” problems in a decomposition-based optimization algorithm to mitigate overfitting in MCEA/D [24]. The algorithm CR-SAEA [25] augments the multiple local classification models approach with a global regression model to help guide its decomposition-based genetic algorithm.
2.2.2 Relationship-Based Surrogates
@@
Modeling the relationships between solutions (e.g. domination) is also an important theme in multi-objective optimization algorithms aided by classification models. The concept of modeling the Pareto-dominance relationship was explored further in some work [26] which compared ten off-the-shelf classification models for use in this task. The algorithm CSEA [27] was introduced which uses a surrogate model to estimate domination between candidate solutions and selected already evaluated reference solutions. The algorithm also uses a measure of uncertainty in the predictions to choose which candidates to evaluate. Two surrogate models: one of the Pareto-domination relationships and a second of the $\theta$-domination relationships between candidates are combined in the algorithm $\theta$-DEA-DP [28]. In REMO [29], the relationships between candidate solutions are modeled with a neural network classifier trained on pairs of solutions which are grouped into equal sized sets based on domination. A related algorithm, named CREMO [30], also explored relationship learning using a convolutional neural network on the decision variables. In FAEA-SDR [31], the strengthened dominance relationship between solutions is modeled instead. Pairwise comparisons of the scalarized objective functions are learned and used in the optimization algorithm PC-SAEA [32]. The concept of domination modeling has also been applied to neural architecture search with promising results [33].
2.2.3 Comparison-Operator Models
@@
Some work has been put into modeling the comparison relationships between the objective functions directly. That is, the value of the comparison operator evaluated on the objective functions between pairs of solutions. The use of these models has remained restricted to the case of single-objective optimization or constraint satisfaction problems. One such study [34] showed an algorithm that ranked solutions using a support vector machine could outperform evolutionary algorithms aided with regression models on some benchmark problems. Binary relationship learning for the pre-selection of solutions was investigated in Hao et al [35]. Multiple classification models were compared for the purpose of ranking solutions based on their comparison relationships in single-objective optimization. A wide variety of pairwise classification models and regression models were compared against each other on benchmark problems for single-objective optimization in [36] with some evidence for the success of the classification approach. A specially designed symmetric neural network architecture was created for the task of modeling the comparison operator between multiple objectives in [37] and applied to solve constraint satisfaction problems in integrated circuit design. While these results demonstrate the potential of a comparison-based model, they did not address the unique challenges of integrating one with a multi-objective optimization algorithm. This includes integrating the comparison predictions across all objectives, diversity preservation in higher dimensions, and dealing with predictions that may not satisfy important symmetries of the comparison operator which are relied on in (for example) domination sorting.
@@
2 $P_{t}\leftarrow$ initial set of $N_{i}$ solutions, Latin hypercube
3 $P\leftarrow$ $N$ solutions selected from $P_{t}$ as in NSGA-II
@@ 11             $P_{m}\leftarrow P_{m}\cup c_{m}$
12             $P_{m}\leftarrow$ select $N$ solutions in $P_{m}$ using model
@@ 14             $w\leftarrow w+1$
      /* Evaluate Solutions */
18       $c\leftarrow$ evaluate $\mu$ random candidate solutions from $P_{m}$
@@ 21       $\theta\leftarrow$ train model on $P_{t}$
      /* Select solutions based on fitness */
@@ 23       $P\leftarrow P\cup c$
24       $P\leftarrow$ select $N$ solutions from $P$ as in NSGA-II
@@
Figure 1: A diagram of the how the CRSEA algorithm (Alg. 1) works. An initial population is generated and the comparison-relationship model trained. Then, the comparison-relationship model is used in the place of the objective functions in an evolutionary algorithm for some number of generations ($w_{\text{max}}$). Several ($\mu$) randomly chosen solutions from the resulting candidate solutions are evaluated and used to retrain the model. The process is repeated until the budget of function evaluations ($\text{FE}_{max}$) is exhausted.
@@ 2.3 Motivation
In addition to the invariance properties of the model, a classification problem is always “simpler” than the equivalent regression problem in the sense that it is solved in the process of solving the regression problem, but does not solve the regression problem itself. That is, a model of the objective function values also gives the classifications whether they predict domination, comparisons, or something else. However, knowledge of what class a solution lies in is not enough information to determine the values of the objective functions. It may be that this “simpler” problem is easier to solve.
The task of learning the comparison relationships between the objectives also forms a naturally balanced classification problem. In a model of the domination relationships, for example, as the algorithm drifts closer to the Pareto front most new examples are dominated by the existing population. The comparisons between objective functions, on the other hand, can always be manipulated to have an equal number of positive and negative examples by swapping the order of the two solutions being compared.
With these arguments and the previously discussed success of comparison-relationship models in solving single-objective optimization problems, we are motivated to use a similar model to drive a genetic algorithm in solving the multi-objective case.
@@ 3 A Comparison-Relationship-Surrogate multi-objective Genetic Algorithm
With an accurate enough model of the comparison relationships (Eq. 1), the expense of evaluating the objective functions in an evolutionary algorithm could be replaced completely with a cheaper call to the model. This approach for integrating information from the model into an optimization algorithm has seen success in other SAEAs (such as in K-RVEA [11]) which motivates us to handle the comparison-relationship model in a similar manner.
With this inspiration, we propose the algorithm described in Alg. 1. The model is first initialized using training points that are drawn from the decision space with Latin hypercube sampling [41] and evaluated using the real objective functions. Candidate solutions are found from the model using an evolutionary algorithm. To combat the model’s loss of accuracy when it is evaluated far from the points it was trained on, the model is periodically updated using a small number of candidate solutions for which the real objective function values are computed. This process (solving the optimization problem and updating the model) is repeated until the budget of function evaluations is exhausted. A diagram of the algorithm is shown in Fig. 1.
The main difference between the new algorithm and existing SAEAs is the use of the comparison-relationship model which directly models the comparison relationships between each objective of every pair of solutions. As discussed in the motivation section, a benefit of this model is that it has the same invariance under monotonic transformations of the objective functions as the genetic algorithms it is being integrated with. It also forms a naturally balanced classification problem solving some difficulties in other classification-based SAEAs. On the other hand, diversity preservation with a model that doesn’t have information about the distance between solutions may be challenging. In the following subsections, we explain in detail the integration of the comparison-relationship model into a genetic algorithm.
3.1 Integrating the Comparison-Relationship Model with an Evolutionary Algorithm
@@
With this restriction in mind, we propose finding optimal solutions from the model using NSGA-II [42]. Briefly summarizing the algorithm, in each generation, a set of children is produced with mutation, crossover, and selection operators (in this paper, polynomial mutation, simulated binary crossover, and tournament selection [43]). After evaluating the objective functions, the population is sorted by increasing rank of non-domination. Within those ranks, the “crowding distance” metric is used to break ties in order to promote diversity. Solutions are selected from the newly sorted list with the smallest rank, largest crowding distance solutions being the most fit. We suggest running the algorithm using only information from the surrogate model.
Although NSGA-II is well suited for use with the comparison-relationship model, there are still a few challenges that come with it. Each is described in detail in the following subsections and we propose ways of mitigating them. First, we find that approximate models of the comparison relationships which do not satisfy the transitive properties must have their results corrected before use in an optimization algorithm. We then discuss the prediction of domination from a model that only predicts strict inequality between solutions. We conclude by talking about the crowding distance metric and the difficulty encountered in using it during selection with a model that only predicts comparisons.
@@ 3.1.1 Violation of the Transitive Property
The comparison operator satisfies the transitive property. For the objective functions (single objective for simplicity) of three solutions, $f_{1}$, $f_{2}$, $f_{3}$, if $f_{1}<f_{2}$ and $f_{2}<f_{3}$ then $f_{1}<f_{3}$ must be true. Approximate models of the comparison operator may violate the transitive property. This can lead to paradoxical situations such as the model predicting a solution 1 dominates solution 2, solution 2 dominates solution 3, but solution 3 is not dominated by solution 1. Such nonsensical results can wreak havoc on algorithms such as the fast non-dominated sorting algorithm from Ref. [42] used in NSGA-II. To integrate a comparison-relationship model into the genetic algorithm, we must find a way of avoiding these situations.
We propose “cleaning” the predictions of the comparison model so that they respect the transitive property. For all solutions $x_{i}$ in $\mathbb{P}$, the population being sorted, we can calculate the noisy predictions from the comparison model. The predictions take the form of a probability that the comparison is true, $\mathcal{P}(f(x_{i})<f(x_{j}))$, where $f(x)$ is the objective function (a single objective for simplicity). To clean the predictions, we calculate the value,
@@
from Ref. [44]. We take the value of $\mathcal{P}(f(x_{i})<f(x_{i}))$ as zero. The quantity $\tau_{i}$ can be interpreted as the probability of the objective function of solution $i$ being smaller than that of a solution chosen at random from $\mathbb{P}$. The comparison predictions defined by $f(x_{i})<f(x_{j}):=\tau_{i}>\tau_{j}$ now satisfy the transitive property. This cleaning procedure is repeated for all objective functions and the domination relationships can be calculated from there, allowing the fast non-dominated sorting algorithm to be run using the comparison model.
3.2 Equality and the Comparison-Relationship Model
Assuming a minimization problem, one solution dominates another if all of its objectives are less than or equal to the other’s and at least one of its objectives is strictly less than the corresponding objective in the other solution. Unfortunately, the comparison-relationship model seems to only predict the case of strict inequality. This prevents us from predicting domination in the cases where some of the objective for two solutions are equal. However, equality happens exactly at the point in between the region where $f(x_{1})<f(x_{2})$ and the region where $f(x_{2})<f(x_{1})$. In terms of the model, these are exactly the points which lie along the classifier’s decision boundary! This allows us to define domination as predicted from the model as $\tau_{i}<=\tau_{j}$ for all objectives and $\tau_{i}<\tau_{j}$ for at least one using our “cleaned” predictions as described in the previous section.